Please show what to write on the blanks. For example (=D7+D8)

Billys Exterminators, Inc., has sales of $734,000, costs of $315,000, depreciation expense of $48,000, interest expense of $35,000, and a tax rate of 35 percent. What is the net income for this firm? Sales Costs Depreciation expense Interest expense $ 734,000 315,000 48,000 35,000 10 35% 12 13 14 15 16 17 18 Tax rate Cash dividends Shares outstanding $ 85,000 110,000
Tax rate Cash dividends Shares outstanding 12 13 35% 85,000 110,000 $ 15 16 17 18 19 20 Complete the following analysis. Do not hard code values in your answers 21 Income Statement S 734,000 315,000 48,000 Sales Costs Depreciation expense EBIT Interest expense EBT Taxes (35%) Net income 23 25 26 27 28 29 30 31 32
Answer #1

EBIT

=(D23-D24-D25)

=Sales-costs-depreciation expense

=734000-315000-48000

=371000

Interest expense=35000

EBT=(D26-D27)

=EBIT-Interest expense

=371000-35000

=336000

Taxes=EBT*Tax rate

=336000*35%=117600

Net income

=(D28-D29)

=EBT-Taxes

=336000-117600

=218400

the above is answer..